If you have a LAN that requires 60 useable hosts, what subnet mask would you assign?
Answer #1

2h - 2 >= required number of hosts

2h - 2 >= 60

h = 6 where h is number of host bits

In total 32 bits host bits = 6 then network bits = 32-6 = 26

In subnet mask network bits are denoted as 1 and host bits are denoted as 0

subnet mask = 11111111. 11111111. 11111111. 11000000

we made 26 network bits as 1 and 6 host bits as 0. Now convert them into decimal

Subnet mask = 255. 255. 255. 192